OpenStudy (anonymous):

The sets of numbers 6, 8, 10 and 5, 12, 13 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples

OpenStudy (jojokiw3):

Do you know the Pythagorean Theorem in equation?

OpenStudy (triciaal):

hint 6,8, 10 same ratio as 3, 4,5

OpenStudy (triciaal):

@Redneckk__Lifee next time do not post a question and leave.

Directrix (directrix):

5, 12, 13 Show that this triple satisfies the Converse of the Pythagorean Theorem: Pythagorean Thereom converse: If the square of one side of a triangle is equal to the sum of the squares of the other two sides, then the triangle is a right triangle. 5^2+ 12^2 = 13^2 25 + 144 = 169 169 = 169 @Redneckk__Lifee I saved the first triple for you.

OpenStudy (dn_20):

Pythagorean Theorem is known as a^2 +b^2 = c^2, and it is only used if and only if the triangle is a right triangle. Some Pythagorean triples, smallest like 3,4,5 , have the same total at the end To met a Pythagorean triples, they must requires: a +b > c and a^2 + b^2 = c^2 Ex: a=3, b=4, c=5 3 +4 >7 \[3^{2} + 4^{2} = 5^{2}\] 9 +16 =25 25 =25 If a +b <c, then it is not a triangle If a +b >c and \[a ^{2} +b ^{2} \neq c ^{2}\] , then it is not a triangle Example: 4,5,6 4 +5 >6 \[4^{2} +5^{2} \neq 6^{2}\]

OpenStudy (anonymous):

okay, we have 6, 8, 10 and 5, 12, 13, as you know the formula is a^2 + b^2=c^2. And now how do we know which one is a,b and c? easy, a and always smaller than c so any 2 number in the given number that smaller than the third, is a and b, the last one is c :) now let's solve it: 6, 8, 10 a^2+b^2=c^2 6^2+8^2=10^2 ( 6 and 8 both smaller than 10 so it makes sense that 6 and 8 is a and b, 10 is c :) ) 36 + 64 =100 100=100 (true ) this would form a right triangle :) 5, 12, 13, 5 and 12 both smaller than 13, so it makes sense that 5 and 12 is a and b, 13 is c :) 5^2 + 12^2 =13^2 25 + 144 = 169 159 = 169. ( false ) a^2+b^2 must =c^2 !!! it can't not be neither bigger or smaller :3 must be equal :3 there you go :3